# Does Welsh Media Need a Review? Detecting Bias in Nation.Cymru’s Political Reporting

### 19 Apr 2026 v1

@@ 1 Introduction
This paper offers the first computational approach to answering this question by examining Nation.Cymru, a prominent and award-winning [^4] Journalists’ Charity, Wales Media Awards 2025, 17 January 2026. Welsh political news outlet. Nation.Cymru describes itself as an independent, national news service that reports on “all political parties without fear or favour.” [^5] Nation.Cymru, “Our only bias is in favour of truth and honesty and against those who lie and deceive”, 6 December 2025. However, its founding charter names Labour, Plaid Cymru, and the Conservatives among parties that “all have something to contribute” [^6] Nation.Cymru, ‘Welcome to Nation.Cymru’, 26 May 2017., notably omitting Reform UK.
@@ 3 Methodology
A key challenge is distinguishing bias about a party from sentiment expressed by a party (e.g., “Plaid criticised the devastating policy” has negative language, but Plaid is the speaker, not the target). I initially tested a dedicated sentiment model but found it lacked the nuanced reasoning needed to distinguish speaker from target.777cardiffnlp/twitter-roberta-base-sentiment-latest, a RoBERTa model fine-tuned on ${\sim}124$M tweets for three-class sentiment. See https://huggingface.co/cardiffnlp/twitter-roberta-base-sentiment-latest. I therefore use the LLM Claude Sonnet 4.6 (Anthropic, 2026) via API as a prompted classifier (prompt available in the code repository). For each biased mention it returns: (1) on_target (1/0): is the bias directed at the named party (off-target mentions are excluded); (2) sentiment score on a 5-point categorical scale $s\in\{-1,-0.5,0,+0.5,+1\}$; (3) one-sentence reasoning. The prompt includes speaker-vs-target examples and a self-check instruction. To validate Stage 2 output, I manually reviewed a stratified random sample of 50 classified mentions and agreed with the LLM’s target-attribution and sentiment labels in 92% of cases; the LLM occasionally incorrectly labeled the on-target party when multiple parties appeared in the same context window.

### 22 Apr 2026 v2

@@ 1 Introduction
This paper offers the first computational approach to answering this question by examining Nation.Cymru, a prominent and award-winning [^4] Journalists’ Charity, Wales Media Awards 2025, 17 January 2026. Welsh political news outlet. Nation.Cymru describes itself as an independent, national news service that reports on “all political parties without fear or favour.” [^5] Nation.Cymru, “Our only bias is in favour of truth and honesty and against those who lie and deceive”, 6 December 2025.
@@ 3 Methodology
A key challenge is distinguishing bias about a party from sentiment expressed by a party (e.g., “Plaid criticised the devastating policy” has negative language, but Plaid is the speaker, not the target). I initially tested a dedicated sentiment model but found it lacked the nuanced reasoning needed to distinguish speaker from target.666cardiffnlp/twitter-roberta-base-sentiment-latest, a RoBERTa model fine-tuned on ${\sim}124$M tweets for three-class sentiment. See https://huggingface.co/cardiffnlp/twitter-roberta-base-sentiment-latest. I therefore use the LLM Claude Sonnet 4.6 (Anthropic, 2026) via API as a prompted classifier (prompt available in the code repository). For each biased mention it returns: (1) on_target (1/0): is the bias directed at the named party (off-target mentions are excluded); (2) sentiment score on a 5-point categorical scale $s\in\{-1,-0.5,0,+0.5,+1\}$; (3) one-sentence reasoning. The prompt includes speaker-vs-target examples and a self-check instruction. To validate Stage 2 output, I manually reviewed a stratified random sample of 50 classified mentions and agreed with the LLM’s target-attribution and sentiment labels in 92% of cases; the LLM occasionally incorrectly labeled the on-target party when multiple parties appeared in the same context window.